OpenStudy (anonymous):

Use the following equation: log2x + log2(x – 7) = 3 To solve the equation, first: Select one: a. Use the quotient property of logarithms to condense the equation. b. Use the product property of logarithms to condense the equation. c. Use the power property of logarithms to condense the equation. d. Either B or C

OpenStudy (anonymous):

@IMStuck

OpenStudy (anonymous):

@rational

OpenStudy (anonymous):

@lncognlto

OpenStudy (lncognlto):

These are the rules of logs: log a + log b = log (a*b) log a - log b = log (a/b) b log a = log (a)^b From this, which do you think would be the appropriate answer?